OpenStudy (pulsified333):

Someone help I will give medal Question involving the rolling of two fair dice. What is the probability of one of the dice showing 3 or the sum being at least 8?

OpenStudy (anonymous):

Well do you know how to compute probability?

OpenStudy (pulsified333):

yes I do, I thought the answer was 16/36 but it was wrong

OpenStudy (anonymous):

So the probability of rolling 3 is equal on /both/ of them so we just leave that probability as 1/6th, not 1/12th. Does that help?

OpenStudy (anonymous):

Huh? On a fair dice there's a 1 in six chance of rolling a 3.

OpenStudy (pulsified333):

yes but can't you roll a 3 on lets say die 1 or die 2?

OpenStudy (anonymous):

Yes, but that's 2/12 chance, which simplifies down to 1/6.

OpenStudy (pulsified333):

oh

OpenStudy (anonymous):

Actually, 16/36 can be simplified. Maybe the answer needs to be simplified. What format is the test?

OpenStudy (pulsified333):

it simplifies it for you

OpenStudy (anonymous):

Oh. well then.

OpenStudy (anonymous):

So we've got 1/6th. Then we need to work out the chance of rolling a sum of of over eight and add it to 1/6th.

OpenStudy (pulsified333):

would it be over 36?

OpenStudy (anonymous):

The denominator?

OpenStudy (pulsified333):

yes

OpenStudy (anonymous):

IDK XD I'll work it out.

OpenStudy (anonymous):

So how many combinations are there that equal above 8?

OpenStudy (pulsified333):

5

OpenStudy (anonymous):

So there is a 5/36 (yeah, it is 36) chance of getting +8. So find a common denominator between 1/6th and 5/36ths and add them together.

OpenStudy (pulsified333):

i made a mistake

OpenStudy (pulsified333):

its the sum of at least 8

OpenStudy (pulsified333):

which equals 15/36

OpenStudy (anonymous):

Ok. So still, multiply 'till you've got a common denominator + add them together.

OpenStudy (pulsified333):

quick question

OpenStudy (anonymous):

Yeah?

OpenStudy (pulsified333):

wouldn't the probability of getting a 3 on one die be 10/36

OpenStudy (anonymous):

I think the probability of 1 showing a 3 is 11/36. http://www.sms.rcs.k12.tn.us/TEACHERS/flaniganj/TEST/chapter_9_files/i0100000.jpg

OpenStudy (anonymous):

Well 36/6 = 6 so it's 6/36

OpenStudy (pulsified333):

where is the 6 coming from?

OpenStudy (pulsified333):

would the answer be 25/36

OpenStudy (anonymous):

You can get a 3 six different way on the first die, and 6 different ways on the second die. But two of those is (3, 3) so you subtract 1 to get 11/36 ways. 11/36 + 15/13 = 26/36

OpenStudy (anonymous):

* 11/36 + 15/36 = 26/36

OpenStudy (pulsified333):

oh i see

OpenStudy (pulsified333):

the answer is still wrong

OpenStudy (anonymous):

Well you're multiplying both sides of the fraction to give it 36 as a denominator, and you need to times 6 by 6 to get to 36, so so you also multiply 1 by six to get from 1/6 to 6/36, which you can add to 15/36.

OpenStudy (anonymous):

Which by my admittedly thirteen-year-old brain is 21/36.

OpenStudy (anonymous):

did you try the 25/36 you came up with? I'm guessing they want you to exclude the double roll.

OpenStudy (pulsified333):

yeah probably because it says one 3 not 2

OpenStudy (anonymous):

was that it?

OpenStudy (pulsified333):

i didn't try it yet. I only have 2 attempts left

OpenStudy (pulsified333):

do you think 25/36 would be the right answer if its not including the double roll

OpenStudy (anonymous):

yeah, not including it, it's 25/36. including it, it's 26/36 or 13/18.

OpenStudy (pulsified333):

no its not it either

OpenStudy (pulsified333):

;(

OpenStudy (anonymous):

red = one 3 blue = sum at least 8